OpenStudy (anonymous):

Write the equation of the line that is perpendicular to a segment created by joining points (1, 0) and (5, -2) and also passes through the midpoint of that segement. Any form is acceptable

OpenStudy (anonymous):

does anyone know how to solve it

OpenStudy (campbell_st):

can you find the slope of the segment..?

OpenStudy (anonymous):

first you have 2 points of a line, so you can get the slope \[\frac{ x _{2}-x }{ y _{2}-y }\] they are perpendicular ,so slope1 * slope 2 = -1 now you got the slope m and get the mid point of the line (1,0(5,-2) and substitute in that form y = mx +c now you are done :d

OpenStudy (campbell_st):

and do you know to find the midpoint

OpenStudy (anonymous):

Midpoint is (3,-1)

OpenStudy (campbell_st):

thats correct... what about the slope of the original segment

OpenStudy (anonymous):

don't know that

OpenStudy (campbell_st):

the line segment has a negative slope... so find the rise and run...  the slope of the segment will be m = - rise/run does that make sense

OpenStudy (anonymous):

so -1/2 @campbell_st

OpenStudy (anonymous):

 and for the second get reciprocal times -1 then you get y= mx+ c and substitute with mid point